# IPPP/25/93 Economical Jet Taggers - Equivariant, Slim, and Quantized Antoine Petitjean1, Tilman Plehn1,2, Jonas Spinner3, and Ullrich Köthe2

### v1

@@ 3.4 Towards online jet tagging
An exciting question is whether small and quantized transformers can meet the resource and latency requirements of FPGAs. In that case, resource-efficient version of the leading jet taggers can be applied at the trigger level. Information about the partonic nature of a jet would open entirely new avenues for HL-LHC analyses. The state-of-the-art approach to FPGA-ready networks is high granularity quantization (HGQ) [Sun:2024soe], where we learn the bit-width of each operation using backpropagation. Since this study can only serve as a first proof of principle, we use GPU implementations instead of an FPGA, as well as our naive fp8 quantization instead of HGQ.
For this test, we reduce a naive transformer and the LLoCa-Transformer to one block and a, typically, 16-dimensional latent space. Even for low numerical precision the orthonormalization has to be done to full precision, so its cost is significant. Therefore, we also show results for the global canonicalization LLoCa-Transformer, which requires the orthonormalization only for one particle and does not involve frame-to-frame conversions. From Fig. 3 we know that L-GATr and L-GATr-slim show poor performance for a minimal number of blocks. Adding more blocks solves this problem, but significantly increases the latency, so we omit the L-GATr architecture for now. The results are shown in Fig. 6. In the left panel we see that, as before, we can reduce the numerical precision without significant performance drop, down to float8.
In the right panel of Fig. 6 we reduce the input data from up to the full set of particle-flow constituents to 30 at most. Depending on the transverse momentum the number of constituents inside a top jet peaks above 50 [Butter:2017cot]. This means that while the general picture does not change, the performance of the top taggers decreases significantly when we only consider the leading 30 constituents.
@@ 4 Outlook
As for industry applications, upscaled transformers are currently driving the performance frontier for jet tagging as the leading ML-application in LHC physics. However, this performance gain comes with an over-proportional compute cost in training and evaluation. For dedicated physics applications, we studied the opposite approach, physics-aware downscaling of transformers, in terms of network parameters and using quantized parameters.
As a starting point, we have presented an efficient version of the equivariant L-GATr architecture, L-GATr-slim. It constructs powerful latent representations for many LHC tasks, from jet tagging to amplitude regression and event generation. Compared to L-GATr, we reduced the training time by a factor six and the memory footprint by a factor 2.3, at no cost in tagging performance. In this way, the geometric algebra approach is competitive with the local canonicalization, specifically the LLoCa-Transformer.
For L-GATr-slim and LLoCa we then tested the performance down to 1k network parameters, either reducing the number of layers or reducing the number of parameters per layer. While this down-scaling comes with a reduced performance, the LLoCa-Transformer remains a powerful tool for a small number of blocks, while L-GATr-slim leads the competition for a fixed number of blocks and a reduced number of parameters per block.
To further improve efficiency, we quantized the network parameters using the standard PARQ and STE methods. The inductive bias from Lorentz equivariance prevents a performance drop from this weight quantization, such that the performance of the LLoCa-Transformer and L-GATr-slim are unaffected by input and by weight quantization. Here, it is crucial that we ensure that the weight quantization does not affect the Lorentz equivariance significantly.
@@ Acknowledgements
We would like to thank Thea Arrestad for expert advice on FPGAs and to Luigi Favaro, Víctor Bresó Pla, and Huilin Qu for many helpful discussions. This work is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant 396021762 – TRR 257 Particle Physics Phenomenology after the Higgs Discovery. It has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101168829, Challenging AI with Challenges from Physics: How to solve fundamental problems in Physics by AI and vice versa (AIPHY). JS is supported by the Humboldt Foundation through a Feodor-Lynen Fellowship. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through the grants INST 35/1597-1 FUGG and INST 39/1232-1 FUGG. This work made use of the facilities of the N8 Centre of Excellence in Computationally Intensive Research (N8 CIR) provided and funded by the N8 research partnership and EPSRC (Grant No. EP/T022167/1). The Centre is coordinated by the Universities of Durham, Manchester and York.
@@ Appendix A Implementation details
Our LLoCa-Transformer closely follows Ref. [Favaro:2025pgz], with a smaller Frames-Net that uses 32 instead of 128 hidden channels in the MLP. We find that this setting achieves the same performance, but reduces FLOPs and memory use significantly.
@@
The L-GATr-slim event generator closely follows the L-GATr event generator in Ref. [Brehmer:2024yqw], which 6 blocks, 32 vector channels, 64 scalar channels, and 8 heads. It uses the same training setup with 200k iterations, AdamW optimizer, batch size 2048, and learning rate , and a step-wise learning rate reduction whenever the loss saturates.
@@
For each network in Fig. 3, we provide the hyperparameters of the architecture in Table 6. The deep models correspond to the right plot. We use different training hyperparameters for each network size, given in Table 7. All networks are trained for 200 000 iterations using the Adam optimizer with a cosine-annealing learning rate schedule. We use the same training hyperparameters for the deep and shallow networks.

### v2

@@ 3.4 Towards online jet tagging
An exciting question is whether small and quantized transformers can meet the resource and latency requirements of FPGAs. In that case, resource-efficient versions of the leading jet taggers can be applied at the trigger level. Information about the partonic nature of a jet would open entirely new avenues for HL-LHC analyses. The state-of-the-art approach to FPGA-ready networks is high granularity quantization (HGQ) [Sun:2024soe], where we learn the bit-width of each operation using backpropagation. Since this study can only serve as a first proof of principle, we use GPU implementations instead of an FPGA, as well as our naive i8+PARQ quantization instead of HGQ.
For this test, we reduce a vanilla transformer and the LLoCa-Transformer to one block and a, typically, 16-dimensional latent space. Even for low numerical precision the orthonormalization has to be done to full precision, so its cost is significant. Therefore, we also show results for the global canonicalization LLoCa-Transformer, which requires the orthonormalization only for one particle, and also does not require frame-to-frame conversions in attention. From Fig. 3 we know that L-GATr and L-GATr-slim show poor performance for a minimal number of blocks. Adding more blocks solves this problem, but significantly increases the latency, so we omit the L-GATr architecture for now.
The results are shown in Fig. 6. In the left panel we see that we can reduce the numerical precision without significant performance drop, down to int8 with static quantization. We find that naively quantizing to ternary weights significantly degrades performance, and conclude that HGQ is necessary to improve further. In the right panel of Fig. 6 we reduce the input data from the full set of particle-flow constituents to the 30 constituents with largest transverse momentum. The number of constituents inside a top jet peaks around 50 [Butter:2017cot]. This means that while the general picture does not change, the performance of the top taggers decreases significantly when we only consider the leading 30 constituents.
@@ 4 Outlook
As for industry applications, upscaled transformers are currently driving the performance frontier for jet tagging as the leading ML-application in LHC physics. However, this performance gain comes with an over-proportional compute cost in training and evaluation. For dedicated physics applications, we studied the opposite approach, physics-aware downscaling of transformers, in terms of reducing network parameters and using quantized parameters.
As a starting point, we have presented an efficient version of the equivariant L-GATr architecture, L-GATr-slim. It constructs powerful latent representations for many LHC tasks, from jet tagging to amplitude regression and event generation. Compared to L-GATr, we reduced the training time by a factor six and the memory footprint by a factor 2, at no cost in tagging performance. In this way, the specialized layer approach is competitive with the local canonicalization approach, specifically the LLoCa-Transformer.
For L-GATr-slim and the LLoCa-Transformer we then tested the performance down to 1k network parameters, either reducing the number of layers or reducing the number of parameters per layer. While this down-scaling comes with a reduced performance, the LLoCa-Transformer remains a powerful tool for a small number of blocks, while L-GATr-slim leads the competition for a fixed number of blocks and a reduced number of parameters per block.
To further improve efficiency, we quantized the network parameters using the standard STE and PARQ methods. The inductive bias from Lorentz equivariance prevents large performance drops from this weight quantization, such that the performance of the LLoCa-Transformer and L-GATr-slim are only mildly affected by input and by weight quantization. Here, it is crucial that we ensure that the weight quantization does not affect the Lorentz equivariance significantly. We have estimated that quantization can reduce the energy consumption by 10.
@@ Acknowledgements
We would like to thank Thea Arrestad for expert advice on FPGAs and to Luigi Favaro, Víctor Bresó Pla, Huilin Qu and Ramon Winterhalder for many helpful discussions. This work is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant 396021762 – TRR 257 Particle Physics Phenomenology after the Higgs Discovery. It has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101168829, Challenging AI with Challenges from Physics: How to solve fundamental problems in Physics by AI and vice versa (AIPHY). JS gratefully acknowledges support from the Alexander-von-Humboldt foundation as a Feodor-Lynen Fellow. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through the grants INST 35/1597-1 FUGG and INST 39/1232-1 FUGG. This work made use of the facilities of the N8 Centre of Excellence in Computationally Intensive Research (N8 CIR) provided and funded by the N8 research partnership and EPSRC (Grant No. EP/T022167/1). The Centre is coordinated by the Universities of Durham, Manchester and York.
@@ Appendix A Implementation details
Our LLoCa-Transformer closely follows Ref. [Favaro:2025pgz], with a smaller Frames-Net that uses 32 instead of 128 hidden channels in the MLP. We find that this setting achieves the same performance, but reduces FLOPs and memory use significantly. Apart from this change, our vanilla transformer and LLoCa-Transformer trainings use the same architecture and training hyperparameters as Ref. [Favaro:2025pgz].
@@
For our ParT trainings, we use the same improved training hyperparameters as for L-GATr-slim, instead of the original training setup [Qu:2022mxj]. We find that these hyperparameters achieve significantly better results, see Tab. 1.
@@
The L-GATr-slim event generator closely follows the L-GATr event generator in Ref. [Brehmer:2024yqw], with 6 blocks, 32 vector channels, 64 scalar channels, and 8 heads. It uses the same training setup with 200k iterations, AdamW optimizer, batch size 2048, and learning rate , and a step-wise learning rate reduction whenever the loss saturates.
@@
For each network in Fig. 3, we provide the hyperparameters of the architecture in Table 7, with ’deep’ networks corresponding to the right panel. We use different training hyperparameters for each network size, given in Table 8. All networks are trained for 200 000 iterations using the Adam optimizer with a cosine-annealing learning rate schedule. We use the same training hyperparameters for the deep and shallow networks.
Quantized taggers
The quantized networks use the same hyperparameters as their non-quantized counterpart. We replace the default two-cycle cosine annealing learning rate schedule for the Transformer and LLoCa-Transformer with a one-cycle cosine annealing schedule when using PARQ. This avoids increasing the learning rate when starting the quantization with the PARQ scheduler. The PARQ scheduler steepness [Jin:2025parq] is set to 100.
For the large taggers, int8 quantization is dynamic for both inputs and weights. For the small taggers, we quantize the weights dynamically during training and we use static quantization for the inputs. The training starts with dynamic quantization, then the quantization parameters are saved after 10 000 steps and updated with each forward pass. To compute and , we use the quantiles and of each input tensor as and , and update the previous values with momentum .